### v1

# C-SEO Bench: Does Conversational SEO Work?

Abstract
Large Language Models (LLMs) are transforming search engines into Conversational Search Engines (CSE). Consequently, Search Engine Optimization (SEO) is being shifted into Conversational Search Engine Optimization (C-SEO). We are beginning to see dedicated C-SEO methods for modifying web documents to increase their visibility in CSE responses. However, they are often tested only for a limited breadth of application domains; we do not understand whether certain C-SEO methods would be effective for a broad range of domains. Moreover, existing evaluations consider only a single-actor scenario where only one web document adopts a C-SEO method; in reality, multiple players are likely to competitively adopt the cutting-edge C-SEO techniques, drawing an analogy from the dynamics we have seen in SEO. We present C-SEO Bench, the first benchmark designed to evaluate C-SEO methods across multiple tasks, domains, and number of actors. We consider two search tasks, question answering and product recommendation, with three domains each. We also formalize a new evaluation protocol with varying adoption rates among involved actors. Our experiments reveal that most current C-SEO methods are largely ineffective, contrary to reported results in the literature. Instead, traditional SEO strategies, those aiming to improve the ranking of the source in the LLM context, are significantly more effective. We also observe that as we increase the number of C-SEO adopters, the overall gains decrease, depicting a congested and zero-sum nature of the problem. Our code and data are available at https://github.com/parameterlab/c-seo-bench and https://huggingface.co/datasets/parameterlab/c-seo-bench.
1 Introduction
The integration of Large Language Models (LLMs) into modern search engines is reshaping how information is delivered to users. Instead of presenting a ranked list of hyperlinks, these systems now generate direct answers in natural language, typically accompanied by inline citations referencing the original web sources. This emerging paradigm, known as Conversational Search Engines (CSEs; Aggarwal et al. 2024), offers a more conversational and informative search experience. Consequently, content providers are increasingly motivated to adopt strategies that improve the visibility and citation frequency and ranking of their web content within CSE-generated answers. Analogous to how Search Engine Optimization (SEO) has long been employed to enhance rankings in traditional search engines, Conversational Search Engine Optimization (C-SEO; Aggarwal et al. 2024) is now emerging to address similar goals in this new context. [^1] https://techcrunch.com/2025/05/08/amazons-newest-ai-tool-is-designed-to-enhance-product-listings/
Current CSE deployments encompass diverse applications with little in common such as product recommendations (e.g., Amazon Rufus and ChatGPT Product Recommendation) and answering questions based on multiple documents (e.g., Perplexity.ai, ChatGPT Search, and Google AI Search). This implies that C-SEO methods for one application or domain may not necessarily work for others. For example, while making a video game description more exciting can be an effective C-SEO method for a CSE-based product recommender, it might not be for a piece of news in a news aggregator. Moreover, due to this increasing adoption of CSE, more actors (e.g., online sellers, or content providers) are concurrently applying C-SEO methods, creating a competitive dynamic similar to that of SEO. Yet, current datasets and benchmarks focus on single applications with only a single actor adopting a C-SEO method (Aggarwal et al., 2024; Pfrommer et al., 2024; Kumar and Lakkaraju, 2024; Nestaas et al., 2025). Thus, we need a comprehensive benchmark covering multiple tasks, domains, and actors to effectively evaluate the domain-specific effects of current C-SEO methods.
Figure 1: C-SEO vs SEO. Comparison of the best C-SEO and SEO methods on our C-SEO Bench across all 6 domains. Left: Best C-SEO strategies still fall behind the best SEO performances. Moreover, C-SEO generally does not introduce any gain (close to 0 boost in ranking). Right: With an increasing rate of actors adopting these methods, actors will experience smaller marginal gain of adopting C-SEO (as in SEO).
In this paper, we introduce a new benchmark, C-SEO Bench, to evaluate C-SEO methods across tasks, domains, and number of actors. In particular, we focus on two common tasks for current CSE applications, answering web questions and product recommendations. In the former, users pose a question with the intent of obtaining an informative answer from multiple web pages, exemplifying online search engines such as Perplexity.ai, ChatGPT Search, or Google AI Search. In the latter, the user asks for a recommendation of a product type and the system responds with a list of products followed by a brief justification. This task represents Amazon Rufus and ChatGPT Product Recommendation. For web questions, we provide three domains covering different applications, such as news, debates, and general web content, while for product recommendations, we provide retail products, video games, and books. We propose to measure the effectiveness of C-SEO methods by the improvement in the citation ranking, i.e., an effective method should lead the LLM to cite the modified document earlier than it cited the original document. This measurement is intuitive, interpretable, and generalizable to any task. Lastly, we propose a new evaluation framework where multiple actors employ C-SEO methods concurrently. In this way, we can analyze the effectiveness of the methods in scenarios with competing actors as in real-life applications.
In our experiments, we show that existing C-SEO methods are not significantly effective for most domains and tasks. Instead, we find that the initial ranking of documents retrieved by the search engine plays a far more dominant role in determining the importance of the documents to the LLM. We also observe that as we increase the number of actors, the overall gains decrease, indicating a congested and zero-sum nature of the problem. These results (illustrated in Figure 1) suggest that traditional SEO methods, those that improve retrieval ranking, remain essential for CSE, in contrary to prior beliefs (Aggarwal et al., 2024). Hence, we believe C-SEO will not replace SEO, but will complement it. The contributions of this paper are as follows:
  • C-SEO Bench. We present the first benchmark to comprehensively evaluate C-SEO methods across multiple tasks, domains, and number of adopting actors.
  • C-SEO is a challenging problem. We provide a comprehensive analysis and comparison of current C-SEO methods and show they largely remain ineffective.
  • SEO remains essential. We show that the ranking of the documents in the LLM context is much more influential than any current C-SEO method to determine the document importance in the LLM response.
2 Background and Related Work
Generating Citations in LLM Responses.
Table 1: Benchmarks comparison. [Uncaptioned image] denotes black-hat C-SEO; [Uncaptioned image] denotes white-hat C-SEO.
Benchmarks Methods Tasks Domains Real Data #Docs #Adopters
Aggarwal et al. 2024 [Uncaptioned image] 1 1 5k 1
Pfrommer et al. 2024 [Uncaptioned image] 1 1 1.1k 1
Kumar and Lakkaraju 2024 [Uncaptioned image] 1 1 10 1
Nestaas et al. 2025 [Uncaptioned image] 1 1 50 Many
C-SEO Bench (Ours) [Uncaptioned image] 2 6 16.3k Many
Generating citations increases LLMs’ trustworthiness by allowing users to verify LLMs’ claims. Therefore, early works on LLMs propose to train models to answer questions citing the documents on which they are based (Nakano et al., 2021; Menick et al., 2022). However, to properly evaluate the quality of the citations, Gao et al. (2023) propose the first benchmark to measure the ability of LLMs to ground and cite sources in their generations. They focus on the fluency, correctness, and citation quality of the responses to verify that the citations are faithful to the source content.
Influencing Citation Ranking.
As with traditional SEO, content providers are incentivized to improve their visibility in conversational search engines. Recent works have explored the possibility of influencing the ranking of a document citation in the response of a language model. While most of them have proposed adversarial methods (i.e., black-hat C-SEO) to evaluate LLM robustness, little work has been done on benign content improvements that make the underlying LLM more convinced about the relevance of the document (i.e., white-hat C-SEO). Table 1 compares our benchmark against current datasets.
[Uncaptioned image]Black-hat C-SEO. Nestaas et al. (2025) introduce preference manipulation attacks, where they add manually-crafted prompt injections to websites to bias LLMs to downgrade a competitor website. Kumar and Lakkaraju (2024) further shows that instead of using manually-crafted prompt injections, it is possible to optimize a string that forces the LLM to start the response with a specific product. Tang et al. (2025) proposes an energy-based optimization method that modifies the target document while avoiding detection. Pfrommer et al. (2024) extends this to black-box models by using a prompting-based jailbreaking attack that generates adversarial instructions. However, most of these works rely on small datasets and on specific domains (see Table 1), limiting their conclusions to the feasibility of these attacks rather than extensively analyzing and measuring their performance. In particular, Nestaas et al. (2025) uses 50 web pages with fictitious cameras, books, and news for their experiments. Kumar and Lakkaraju (2024) base their experiments on the descriptions of ten fictitious coffee machines and the promotion of two of them. Only Pfrommer et al. (2024) conduct their experiments on a larger dataset of 1k products, however, of only five categories.
[Uncaptioned image]White-hat C-SEO. Aggarwal et al. (2024) propose and evaluate multiple white-hat C-SEO methods based on stylistic changes of web content so that it boosts its visibility, measured as word count, on the LLM generation. They released a dataset of 1k information-seeking queries with five websites paired to each query. However, this benchmark only focuses on a single task, answering web questions, and does not cover diverse domains. These constraints limit the analysis of the domain-specific effects. Moreover, the evaluation is based on the word count, i.e., the number of words used by the LLM to discuss a document. This metric does not reflect improvements in tasks such as product recommendation. Lastly, their evaluation is limited to scenarios where only a single actor adopts C-SEO methods. Consequently, all these limitations prevent the evaluation of C-SEO methods in realistic situations. Lastly, Bardas et al. (2025) propose another prompting method to improve the relevance of a document for a specific query. However, their method assumes knowing the user query beforehand, which limits its applicability to real scenarios, where user queries are unknown. In our work, we aim to evaluate content improvements rather than the robustness of LLMs, so we employ white-hat C-SEO methods.
3 Problem Definition
Figure 2: Conversational search engine setup. We illustrate the data pipeline for product recommendation. After applying a C-SEO method on the third document, its ranking gets boosted by $+2$ positions.
In this section, we formalize the task of Conversational Search Engine Optimization (C-SEO), illustrated in Figure 2. In a Conversational Search Engine (CSE), a language model conditions its outputs on a set of external documents retrieved by a traditional search engine to deliver the information in a conversational way and providing inline citations referring to the original sources.
Let $q\in\mathcal{Q}$ denote a query, and $\mathcal{D}_{q}=\{d_{1},d_{2},\dots,d_{n}\}$ a set of relevant documents retrieved by a search engine and used as context for the LLM $\phi$. The LLM samples an output conditioned on the query and documents in context:
$(t_{1},t_{2},\dots,t_{k},\ [c_{1}],t_{k+1},\dots,t_{m},\ [c_{2}],\dots,\ [c_{n}])\sim\phi(q,\mathcal{D}_{q})$ (1)
where each $C_{i}$ is the index of a $d\in\mathcal{D}_{q}$.
C-SEO is the task of modifying a document $d_{i}$ such that it improves its position in the sequence of citations $c_{1},c_{2},\dots,c_{n}$, i.e., $d_{i}$ is cited earlier by $\phi(q,\mathcal{D})$, with $C_{1}$ being the citation of $d_{i}$ as best case scenario. Formally, let $f(q,d_{i})$ be a scoring function implicitly learned by the LLM that governs the relevance or likelihood of citing $d_{i}$ in response to query $q$. The goal of C-SEO is to modify $d_{i}$ to maximize this score: $\max_{d_{i}}f(q,d_{i})$. Equivalently, we aim to minimize the citation rank, $\text{rank}(c_{i})$, defined as the position of $d_{i}$ in the ordered citation list of the LLM output: $\min_{d_{i}}\text{rank}(c_{i})$.
CSE systems usually consist of two steps: retrieval and generation. Accordingly, we define two families of methods aimed at improving citation rankings: (i) Search Engine Optimization (SEO), which focuses on improving a document’s position in the retrieval ranking, with the expectation that this will indirectly influence the citation ranking of the LLM output ; (ii) C-SEO that improves the document to make it more appealing to the LLM.
Multiple C-SEO Adopters Setup.
We model the interaction among content providers as a non-cooperative game. Each player $i$ controls a document $d_{i}$ and aims to maximize their utility $u_{i}$, defined in terms of their citation rank. The strategy space $\mathcal{S}_{i}$ includes all C-SEO methods applicable $d_{i}$. To better understand the method’s performance under different competitive assumptions, we introduce the notion of adoption rate. The adoption rate specifies the percentage of players, $0\leq\alpha\leq 1$, adopting a method.. Unilateral adoption is a special case, where a single player applies a method to its document while all others remain static. Most prior works have focused on this scenario (Kumar and Lakkaraju, 2024; Tang et al., 2025; Pfrommer et al., 2024; Aggarwal et al., 2024).
4 C-SEO Bench
C-SEO Bench is a benchmark designed to evaluate C-SEO methods across two tasks: product recommendation and question answering. Each task spans multiple domains to capture various real-world use cases and evaluate domain-specific effects.
Table 2: Benchmark statistics.
Task Domain #Queries #Docs
Product Recommendation Retail 500 5000
Video Games 436 4360
Books 249 2245
Question Answering Web 300 1500
News 294 2375
Debate 142 880
Product Recommendation.
This task involves ranking items based on their relevance to a user’s query, with the requirement that each position in ranking must be justified. We assume a cold start setting, where no information about the user is available. The user submits a query, and the system retrieves 10 documents (product descriptions), which are presented in an unsorted order. The LLM must then recommend the best five products based on the provided descriptions, using only the content of the retrieved documents. The domains in this task include: (i) Retail: General e-commerce products from the Amazon website, (ii) Video games: video game descriptions from the Steam platform, and (iii) Books: Book descriptions from Google Books API.
Question Answering.
This task evaluates the ability to answer questions based on a list of documents. The domains in this task include: (i) Web Questions: real user queries on Google Search, (ii) News: questions about a world event covered by multiple news pieces, and (iii) Debate: Opinionated or controversial questions drawing on content representing diverse viewpoints.
With these tasks and domains, we aim to cover the most prominent use cases of CSE applications. Unlike prior works (Nestaas et al., 2025; Kumar and Lakkaraju, 2024; Pfrommer et al., 2024; Aggarwal et al., 2024), our benchmark allows the development of methods tailored to specific domains and the study of their generalizability.
4.1 Datasets Construction
Retail.
We build our retail dataset on the Amazon Shopping Queries dataset (Reddy et al., 2022). We select English-language queries and include products labeled as either relevant (E) or somewhat relevant (S) to those queries. For each product, we construct a document by concatenating its bullet point description and its product description. The resulting dataset contains 500 unique queries. Each query is associated with 10 relevant products, each described by its corresponding product document.
Video Games.
We build our video games dataset based on the Steam Games dataset.222huggingface.co/datasets/FronkonGames/steam-games-dataset We treat the Search Tag field as a proxy for user queries. We use the short description of the Steam platform game page (About the Game) as the product description. The resulting dataset contains 436 unique queries, with 10 games per query.
Books.
For the books dataset, we generate queries using GPT-4o by first asking for the most popular book genres, and then requesting short queries (maximum of three words) that correspond to searches for books by mood or theme within those genres (prompt in Appendix F) We retrieve book titles and descriptions from the Google Books API. As part of the data cleaning process, we remove book descriptions shorter than 200 characters and discard any queries with fewer than seven associated books. The resulting dataset has 249 queries with nine documents per query on average.
Web.
For the web dataset, we sample 300 questions from the development set of Kwiatkowski et al. (2019). We generate three variations of every question using query expansion with GPT-4o-mini (prompt in Appendix F). Then, we use each expanded query to search the web using the Brave Search API and retrieve the most relevant pages, collecting five pages per original question. Lastly, we extract the main textual content from these pages using the boilerpy3 library, which strips boilerplate and HTML to retain meaningful content snippets.
News.
We build our news dataset based on the Multi-News dataset (Fabbri et al., 2019). Each example of this dataset consists of a set of news articles reporting on the same real-world event, and a summary synthesizing the main points of the associated news articles. We select instances with at least seven news articles and generate a single question for each instance using GPT-4o-mini (prompt in Appendix F), conditioned on the available summary. We clean each document by removing newline characters and duplicated spaces, and by filtering out documents with fewer than 50 words to avoid low-quality or incomplete content. The resulting dataset has 294 queries, with eight documents per query on average.
Debate.
We build our debate dataset using the data from Liu et al. (2023), which contains queries on various topics along with a list of text snippets from multiple websites presenting both supporting and opposing viewpoints. While the original task was to verify statements based on the snippets, we reuse the original queries and their accompanying snippets as documents. We remove the queries with less than five associated snippets.
These six datasets compose our C-SEO benchmark, which then covers six domains grouped in two common CSE tasks, question answering and product recommendation. In total, our C-SEO benchmark contains more than 1.9k queries and 16k documents. This setup allows us to assess the generalization and domain-specific effects of C-SEO methods.
4.2 Evaluation
Let $\mathcal{D}_{q}$ be the set of documents retrieved for query $q$ with $|\mathcal{D_{\text{q}}}|=n$. Let $\mathcal{Q}$ be the set of all queries in a domain of our benchmark, $\alpha\in[0,1]$ the proportion of documents (i.e., players) adopting a C-SEO method, and $\mathcal{D}_{\text{q}}^{\alpha}$ the subset of $\mathcal{D}_{q}$ implementing a C-SEO method. For each document $d\in\mathcal{D}_{\text{q}}^{\alpha}$, we extract the position of the document citation in Equation 1 and call it rank. We define the rank improvement (i.e., gain) for the document $i$ for query $q$ as: $\Delta_{q,d}=\text{rank}_{d}^{0}-\text{rank}_{d}^{\alpha}$. $\text{rank}_{d}^{0}$ is the baseline citation rank (i.e., no C-SEO methods are applied) and $\text{rank}_{d}^{\alpha}$ is the citation rank after C-SEO optimization (at adoption rate $\alpha$). The average gain for a C-SEO method at adoption rate $\alpha$ is: $\Delta(\alpha)=\frac{1}{|Q|}\sum_{q\in Q}\frac{1}{|\mathcal{Q_{\text{q}}^{\alpha}}|}\sum_{d\in\mathcal{D_{\text{q}}^{\alpha}}}\Delta_{q,d}^{\alpha}$.
We access the statistical significance of the rank improvement using the right-tailed Wilcoxon signed-rank test. This test allows us to evaluate whether the rank after applying a C-SEO method is statistically significantly smaller than the original rank. We adjust the p-values using the Holm-Bonferroni correction (Holm, 1979) to prevent the multiple comparison problem and to be able to compare multiple C-SEO methods within each domain.
[…]
5 C-SEO Methods
Our main research question is whether C-SEO methods can increase the citation ranking of a document in a conversational search engine for any task and domain. To investigate this, we benchmark the following content transformation C-SEO methods introduced by Aggarwal et al. (2024):
  1. 1.
    Authoritative: Modifies the text to enhance its authority and persuasiveness.
  2. 2.
    Statistics: Introduces statistical elements to increase the perceived technical depth.
  3. 3.
    Citations: Adds references to increase credibility and trustworthiness.
  4. 4.
    Fluency: Polishes the text to improve grammar and coherence.
  5. 5.
    Unique Words: Incorporates less common words to increase the perceived uniqueness of the text.
  6. 6.
    Simple Language: Simplifies the documents for an easier reading.
  7. 7.
    Quotes: Adds quotations to increase trust in the document.
In addition, we propose to benchmark two new C-SEO methods:
  1. 1.
    Content Improvement: The combination of all the seven transformations above into a single holistic text improvement method. An example is shown in Figure 3.
  2. 2.
    LLM Guidance: Inspired by the LLMs.txt standard333llmstxt.org, this novel prompt-based method generates a markdown summary that is concatenated at the beginning of the original document to guide the LLM about its content.
Overall, we benchmark a total of nine methods. Appendix E details the prompts used by each method. These methods are used to improve the documents in $\mathcal{D_{\text{q}}}$ defined in Section 3. Then, we input the user queries with their lists of documents $\mathcal{D_{\text{q}}}$ to an LLM to generate the answers citing those documents, as in eq. 1. We use gpt-4o-mini-2024-07-18 to run all the C-SEO methods. We run the Conversational Search Engine with gpt-4o-mini-2024-07-18 and claude-3-5-haiku-20241022.
Figure 3: Example of C-SEO transformation. Content Improvement makes the text more attractive by highlighting key features (e.g., by bolding) and structuring the text to make the information more accesible.
6 Experiments
6.1 Main Results
Table 3: C-SEO Bench results. Average and standard deviation of the rank improvements for GPT-4o-mini. Bold values are statistically significant ($p<0.05$) using Bonferroni-Holm correction.
Product Recommendation Question Answering
Method Retail Games Books Web News Debate
Authoritative 0.11 ±1.18 0.07 ±1.25 0.11 ±0.94 -0.04 ±0.89 -0.04 ±1.03 0.01 ±1.55
Statistics -0.07 ±1.00 0.00 ±1.11 -0.11 ±1.04 -0.53 ±1.27 -0.05 ±1.31 -0.80 ±1.70
Citations -0.01 ±1.14 0.04 ±1.23 0.00 ±0.91 0.03 ±1.00 -0.10 ±1.09 -0.15 ±1.50
Fluency 0.06 ±1.11 0.07 ±1.27 0.07 ±1.05 0.03 ±0.89 -0.01 ±1.08 0.32 ±1.64
Unique Words 0.09 ±1.09 0.02 ±1.20 0.04 ±1.04 -0.07 ±0.92 -0.10 ±1.03 -0.08 ±1.66
Tech. Terms 0.05 ±1.13 0.07 ±1.32 -0.03 ±0.86 0.01 ±0.95 -0.03 ±0.97 -0.05 ±1.56
Simple Lang. 0.03 ±1.05 0.11 ±1.39 0.01 ±0.78 0.02 ±1.02 -0.04 ±1.00 0.04 ±1.63
Quotes 0.06 ±1.09 0.04 ±1.29 0.01 ±0.97 0.00 ±0.96 -0.06 ±1.07 -0.19 ±1.64
LLM Guid. 0.36 ±1.47 0.24 ±1.05 0.14 ±1.07 0.10 ±0.99 0.00 ±1.10 0.15 ±1.65
Content Impr. 0.18 ±1.09 0.13 ±1.23 0.11 ±0.95 0.02 ±0.90 -0.04 ±1.00 0.11 ±1.61
Best SEO 2.77 ±2.31 1.89 ±2.32 1.60 ±2.04 0.87 ±1.35 0.70 ±1.64 1.54 ±2.07
Table 3 shows that most existing C-SEO methods do not achieve significant gains for unilateral adopters across tasks or domains. In many cases, these methods leave the rankings unchanged. For example, 61.0% of the retail product ranks are the same after applying the LLM guidance. When changes do occur, they tend to cause relatively large shifts. But, these shifts are both positive and negative, and partially cancel each other out. For example, the LLM guidance on the retail data has a positive boost in 26.2% of the cases and a negative boost in 12.8% of the cases. As a result, the overall average effect is close to zero, but with high variance.
Out of 54 cases, we uncover only three where the ranking improvements are statistically significant. As detailed in section 4.2, we apply the Wilcoxon signed-rank test to check the statistical significance of positive ranking boosts. Among all evaluated methods, only LLM guidance and content improvement yield statistically significant gains. But these gains are only significant on the retail domain for content improvement and only on the retail and video games domains for LLM guidance. We found no method effective for the question answering task, and no methods effective for the Haiku 3.5 model (results in appendix). These findings highlight the very limited effectiveness of most stylistic or content-editing strategies: when a C-SEO method significantly boosts the ranking, it does so only for specific domains, tasks, and models.
LLM guidance shows promising results in some specific cases. While its effectiveness is not consistent across all settings, its relative success suggests a potential direction for future work: rather than focusing solely on rewriting and replacing content for LLMs, it may be beneficial to augment content in ways that enhance both machine interpretability and human readability.
While the results from Aggarwal et al. (2024) show some initial optimism about C-SEO methods, we do not observe the same effectiveness. The differences are due to the metrics used. They report their main results as word count, i.e., the ratio of words used in the LLM response to talk about the target document. However, this metric does not measure the LLM preference, contrary to the citation ranking that we use. A higher word count does not necessarily correspond to a better citation ranking. Furthermore, a careful inspection of their results confirms ours. They also show the results with a position-adjusted word count to consider citation rankings. However, their results with this metric show a general decrease in the scores, implicitly indicating that the C-SEO methods do not generally improve citation ranking. Therefore, the results of both papers on LLM preferences do not contradict each other.
6.2 Importance of Traditional SEO
While C-SEO methods aim to modify documents to condition the LLM to mention them earlier in their response, traditional SEO tries to improve the ranking of a document in the retrieval component and consequently in the LLM context. Figure 5 shows the ranking improvement of documents after positioning them in the $i^{th}$ position in the LLM context. As expected, the top-3 positions lead to the largest gains. Since target documents are selected randomly, the last three positions (i.e., 8, 9, and 10) lead to performance degradation because many documents would be moved to worse positions than their original ones.
Comparing this figure with Table 3 shows that making the target document the first one in the LLM context window leads to far greater citation ranking gains in the LLM response than any C-SEO method. This indicates that the position of a document in the input context has a dominant effect on how the LLM ranks or selects content. This finding challenges the core assumptions of Aggarwal et al. (2024) that traditional SEO will be outdated by C-SEO. Our results show that sophisticated content modifications are often outweighed by improvements in document ordering in the LLM context. As a result, traditional SEO strategies remain critical for the visibility of content creators in conversational search engines. Therefore, C-SEO methods must be considered as a complement and not a replacement for traditional SEO.
Figure 4: Importance of traditional SEO. Average boost in ranking (y-axis) when the document is placed at a specific position in the LLM context (x-axis, SEO Baseline). The boosts of (at least) the first two positions are significant for all domains
Figure 5: C-SEO is a zero-sum game. The average gain per adopter decreases (y-axis) with the increasing number of adopters (x-axis), for retail (dashed) and video games (dotted) on gpt-4o-mini.
6.3 C-SEO Behaves as a Zero-Sum Game
We simulate C-SEO adoption as a competitive game between content creators to better understand how these methods behave under partial or full adoption. Figure 5 reports the area under the curve (AUC) metric, defined in Section 4.2, for the best two C-SEO methods (i.e., LLM guidance and content improvement) in the two domains with the highest gains (i.e., retail and video games). The AUC confirms that the LLM guidance provides the highest ranking boost on average for varying adoption rates. So, even when the adoption rate is unknown, LLM guidance is more likely to improve ranking than content improvement. Moreover, Figure 5 shows the average ranking improvements of both C-SEO methods as a function of the rate of adoption. We observe that while early adopters experience significant gains, these gains decrease steadily as more players adopt the same method. This pattern reveals two key characteristics of C-SEO: it is both a congested and a zero-sum game.
The congestion effect arises because the benefits of adopting an effective method are diluted as more adopters do the same. Improvements in one citation’s rank necessarily come at the expense of others, creating competitive pressure and the need to continuously create new and better methods. As a result, while C-SEO methods benefit early adopters, their advantage decreases over time, eventually converging toward zero at full adoption. These results for white-hat C-SEO methods complement the results of Nestaas et al. (2025): they find that under the black-hat/adversarial setting, the C-SEO game leads to a prisoners’ dilemma. Different setups lead to different dynamics of real-world adoption. Overall, our results highlight the need to evaluate C-SEO methods not only by their unilateral adoption gains but also by their dynamics from partial to full adoption.
7 Conclusion
This work introduces a new benchmark to evaluate conversational search engine optimization methods (C-SEO), spanning two tasks and six domains that reflect the most prevalent real-world use cases of conversational search engines. Our analysis reveals that current C-SEO methods are largely ineffective in influencing citation ranking. However, traditional SEO methods, those aiming to improve the retrieval ranking, have a significantly larger impact on the citation rankings in the LLM output. This result shows the importance of traditional SEO in the new conversational search engine applications. By analyzing the problem from a game theory perspective, we have shown that C-SEO is a congested, zero-sum game. This highlights the incentives for content creators to keep investing in new methods to maintain a competitive advantage.
Limitations and Future Work.
Despite the initial excitement for C-SEO methods that condition LLMs to increase the likelihood of citing a document first, our experiments show that they remain largely ineffective. Nevertheless, our conclusions apply exclusively to white-hat C-SEO methods, leaving black-hat C-SEO methods outside our scope. We do not aim to evaluate the robustness of LLMs against malicious content manipulation because it is a different problem that needs an adaptive adversarial evaluation (Carlini et al., 2019; Tramer et al., 2020), beyond our white-hat C-SEO scope. Moreover, we do not examine the potential interplay between traditional SEO and C-SEO methods. Future work could explore new text modification strategies that explicitly target both SEO and C-SEO to explore possible synergistic or antagonistic interactions. Our experiments are conducted with commercial proprietary models. Lastly, our dataset is restricted to English-language content, which may reduce the generalizability of our findings to other linguistic or cultural contexts.
Acknowledgements
This work was supported by the NAVER corporation.
Appendix A Broader Impacts
Our work analyzes whether current white-hat conversational search engine optimization methods can improve the visibility of online documents such as websites or product descriptions in the responses on conversational search engines. These white-hat techniques are ethical in that they modify the content of the only documents that the content creator owns. Although, prior works have shown the existence of black-hat methods, i.e., adversarial attacks that downgrade the citation ranking of competitors. Our benchmark is not designed for black-hat methods, but could potentially be used to test these attacks. This is a dual-use: adversarial evaluation informs researchers and the public of the existence of potential vulnerabilities in an existing system, but can also be used by malicious actors.
Appendix B Experimental Setup
We use gpt-4o-mini-2024-07-18 to run all the C-SEO methods and run the Conversational Search Engine with gpt-4o-mini-2024-07-18 and claude-3-5-haiku-20241022. We run the Conversational Search Engine with gpt-4o-mini-2024-07-18 and claude-3-5-haiku-20241022. We use the default decoding parameters from the model providers. In the experiments with increasing number of actors, we use use a maximum of ten. So 10% means one actor, and 100% means ten.
Appendix C License and Terms of Use of Used Assets
We used the following datasets with the corresponding licenses. In all cases, they allow the use of their data for research purposes, so we agree to their terms of use.
  • Amazon Shopping Queries [Reddy et al., 2022]. License: Apache 2.0
  • Steam Games from FronkonGames. License: CC-by-4.0.
  • Natural Questions [Kwiatkowski et al., 2019]. License: CC-by-3.0.
  • Multi News [Fabbri et al., 2019]. They use a custom license allowing the use for non-commerical and research purpouses.
  • Debate [Liu et al., 2023]. License: MIT.
Appendix D Other Results
Table 4: C-SEO Bench results on the Haiku 3.5 model. None of the LLM Document Optimization methods achieve significant gains. Best SEO is always statistically significant.
Product Recommendation Question Answering
Method Retail Games Books Web News Debate
Authoritative -0.53 ±1.29 -0.20 ±1.05 -0.34 ±1.06 0.03 ±0.72 0.04 ±0.77 0.01 ±1.41
Statistics -0.82 ±1.47 -0.10 ±1.01 -0.60 ±1.34 -0.58 ±1.18 -0.12 ±0.91 -0.85 ±1.66
Citations -0.49 ±1.31 -0.16 ±0.96 -0.31 ±1.10 -0.06 ±0.78 -0.09 ±0.87 0.06 ±1.31
Fluency -0.48 ±1.30 -0.08 ±1.08 -0.37 ±1.18 0.00 ±0.73 0.05 ±0.87 0.18 ±1.40
UniqueWords -0.52 ±1.26 -0.22 ±0.91 -0.43 ±1.11 -0.01 ±0.83 -0.05 ±0.87 0.04 ±1.40
Tech. Terms -0.45 ±1.26 -0.19 ±0.98 -0.39 ±1.17 -0.04 ±0.90 -0.07 ±0.81 -0.01 ±1.53
Simple Lang. -0.43 ±1.31 -0.09 ±0.96 -0.32 ±1.18 -0.01 ±0.68 0.01 ±0.84 0.01 ±1.48
Quotes -0.50 ±1.27 -0.14 ±1.07 -0.33 ±1.18 -0.08 ±0.93 -0.03 ±0.76 0.03 ±1.33
LLM Guid. -0.32 ±1.36 0.00 ±1.13 0.06 ±1.21 0.02 ±0.83 -0.09 ±0.97 0.12 ±1.60
Content Impr. -0.29 ±1.28 -0.13 ±1.02 -0.08 ±1.08 -0.13 ±0.96 -0.09 ±0.91 0.11 ±1.48
Best SEO 1.61 ±1.96 0.93 ±1.59 0.61 ±1.48 0.35 ±1.22 0.31 ±1.37 0.56 ±1.81
Table 5: P-Values for Haiku 3.5. Complementary table to Table 4.
Product Recommendation Question Answering
Method Retail Games Books Web News Debate
Authoritative 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
Statistics 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
Citations 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
Fluency 1.0000 1.0000 1.0000 1.0000 1.0000 0.8276
Unique Words 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
Technical Terms 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
Simple Lang. 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
Quotes 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
LLM Guid. 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
Content Impr. 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
Best SEO 0.0000 0.0000 0.0000 0.0000 0.0007 0.0050
Table 6: SEO Baseline results Average and standard deviation of the rank improvements for gpt-4o-mini according to the document position in the LLM context. This table contain the same data as Figure 5.
Product Recommendation Question Answering
In-context order Retail Games Books Web News Debate
1 2.77 ±2.31 1.89 ±2.32 1.60 ±2.04 0.87 ±1.35 0.70 ±1.64 1.54 ±2.07
2 1.78 ±2.08 1.28 ±2.01 1.28 ±1.95 0.19 ±1.31 0.45 ±1.51 0.41 ±1.81
3 0.67 ±2.11 0.57 ±1.85 0.48 ±1.65 -0.22 ±1.09 -0.01 ±1.44 -0.37 ±1.80
4 0.06 ±2.02 0.13 ±1.74 0.03 ±1.63 -0.30 ±1.20 -0.13 ±1.46 -0.36 ±1.71
5 -0.18 ±2.06 -0.12 ±1.73 -0.27 ±1.61 -0.53 ±1.23 -0.21 ±1.48 -0.53 ±1.93
6 -0.47 ±1.99 -0.39 ±1.75 -0.47 ±1.75 -1.74 ±1.64 -0.17 ±1.50 -1.06 ±2.07
7 -0.67 ±1.96 -0.48 ±1.68 -0.50 ±1.62 -1.79 ±1.62 -0.50 ±1.50 -1.73 ±1.91
8 -0.76 ±1.92 -0.58 ±1.67 -0.59 ±1.67 -1.79 ±1.64 -0.70 ±1.56 -2.08 ±1.72
9 -0.80 ±1.91 -0.58 ±1.63 -0.88 ±1.75 -1.74 ±1.60 -1.15 ±1.77 -2.14 ±1.79
10 -0.78 ±1.93 -0.64 ±1.70 -1.02 ±1.68 -1.79 ±1.66 -1.29 ±1.68 -2.18 ±1.66
Table 7: P-Values from the SEO baselines across all domains. Complementary table to Table 6.
Product Recommendation Question Answering
In-context order Retail Games Books Web News Debate
1 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
2 0.0000 0.0000 0.0000 0.0344 0.0000 0.0634
3 0.0000 0.0000 0.0000 1.0000 1.0000 1.0000
4 1.0000 0.3271 1.0000 1.0000 1.0000 1.0000
5 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
6 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
7 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
8 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
9 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
10 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
Table 8: P-values for C-SEO methods on gpt-4o-mini. Complementary table to Table 3.
Product Recommendation Question Answering
Method Retail Games Books Web News Debate
Authoritative 0.1981 0.7140 0.3377 1.0000 1.0000 1.0000
Statistics 1.0000 0.9056 1.0000 1.0000 1.0000 1.0000
Citations 1.0000 0.9056 1.0000 1.0000 1.0000 1.0000
Fluency 0.4738 0.7140 0.6085 1.0000 1.0000 0.1633
Unique Words 0.2674 0.9056 1.0000 1.0000 1.0000 1.0000
Technical Terms 0.6428 0.7786 1.0000 1.0000 1.0000 1.0000
Simple Lang. 0.8286 0.5475 1.0000 1.0000 1.0000 1.0000
Quotes 0.6016 0.8998 1.0000 1.0000 1.0000 1.0000
LLM Guid. 0.0000 0.0000 0.1320 0.3069 1.0000 1.0000
Content Impr. 0.0008 0.1373 0.2078 1.0000 1.0000 1.0000
Best SEO 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
Appendix E Prompts for C-SEO Methods
All methods from [Aggarwal et al., 2024] used the same prompt as the original paper and we refer the reader to them. The prompts for our methods are shown below.
Prompt for LLM Guidance
Create a llms.txt markdown file to provide LLM-friendly content.
This file summarizes the main text and offers brief background
information, guidance, and links (if available).
Follow this template
# Title
> Introduction paragraph
Optional details go here
## Section name
More details
Here is the content of the text:
{text}
Prompt for Content Improvement
Rewrite the following text to make it more fluent, authoritative, and
persuasive without altering the core content. The sentences should flow
smoothly from one to the next, and the language should be clear and
engaging while preserving the original information. The revised text
should reflect confidence, expertise, and assertiveness while
maintaining the original content’s meaning and relevance. The text
should be assertive in its statements, such that the reader believes
that this is a more valuable source of information than other texts.
Lastly, give structure to the text.
{text}
Appendix F Prompts for Dataset Construction
We used two prompts to generate the dataset of book queries. The first prompt was used to generate a list of 500 book queries. However, it did not generate 500 queries. Therefore, we used the second prompt to generate the remaining queries.
Prompt for Book Queries
generate 500 book queries by genre, mood, theme. Try to be as diverse as
possible. Return them in a python list. make the queries short, like 3
words max
Second Prompt for Book Queries
User: what are the most selling book genres?
Assistant: ...
User: generate 10 search queries for mood and theme of books with max 3
words for each sub-genre you listed. return it as a json
Query Expansion for Web Dataset
You are an expert at generating search queries for a search engine.
Generate {num_queries} search queries that are relevant to this
question. Output only valid JSON.
User question: {query}
Format: {"queries": ["query_1", "query_2", "query_3", "query_4"]}
Query Generation for News
Generate 1 question for the following piece of news article:
{article}. You should return a json with the key ’questions’ and
a list of questions as the value.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our main claims are summarize in Figure 1 and Section 1. Section 6 provide evidence for them.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Discussed in Section 7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This is not a theoretical paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Appendix B explains the setup and the hyperparameters. We also release our code and data with the submission.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Code and data are available in the Github and Huggingface links provided in the abstract.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Appendix B provides the details of how we run the models. Prompts are provided in Appendix E.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Section 4.2 explains how we conduct the statistical significance of the experiments. Table 3 and Table 4 provides means, standard deviations, and statistically significant results in bold.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [N/A]
  39. Justification: We used commercial APIs to run the models.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have read and understood the code of ethics. We have done our best to conform.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Discussed in Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We use publicly available models. Our data do not pose risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Discussed in Appendix C.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our data has a dataset card in Huggingface and our code has a README.md with enough explanations to run it. The creation of the dataset is explained in Section 4.1 Used prompts are provided in Appendix F.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We did not use crowdsourcing nor experiments with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not conduct experiments with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLM assistants were only used for spell checkers, grammar suggestions, and programming aid for editing purpouses.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 a list of questions as the value.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: Our main claims are summarize in Figure 1 and Section 1. Section 6 provide evidence for them.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: Discussed in Section 7.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: This is not a theoretical paper.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: Appendix B explains the setup and the hyperparameters. We also release our code and data with the submission.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: Code and data are available in the Github and Huggingface links provided in the abstract.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: Appendix B provides the details of how we run the models. Prompts are provided in Appendix E.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34.
-Justification: Section 4.2 explains how we conduct the statistical significance of the experiments.
-Table 3 and Table 4 provides means, standard deviations, and statistically significant results in bold.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [N/A]
-  39. Justification: We used commercial APIs to run the models.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We have read and understood the code of ethics. We have done our best to conform.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: Discussed in Appendix A.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: We use publicly available models. Our data do not pose risk of misuse.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: Discussed in Appendix C.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64.
-Justification: Our data has a dataset card in Huggingface and our code has a README.md with enough explanations to run it.
-The creation of the dataset is explained in Section 4.1 Used prompts are provided in Appendix F.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: We did not use crowdsourcing nor experiments with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: We do not conduct experiments with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: LLM assistants were only used for spell checkers, grammar suggestions, and programming aid for editing purpouses.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.